Given the IEEE-754 Single Precision Floating Point number (stored Excess-127)111111000101011000000000000000002, determine the following: a.The sign of the mantissa b.The magnitude of the exponent (convert to base 10) c.The sign of the exponent
Answer #1

IEEE-754 single precision floating pointing number representation:

Sign (S) Exponent (E) Mantissa (M)
1 bit 8 bits 23 bits

Given Floating Point number = 2

Sign Exponent Mantissa

Implicit Normal form:

Value = (-1)S (1.M) x 2E-127

M = 101011000000000000000002

M = 1 x 2-1 + 0 x 2-2 + 1 x 2-3 + 0 x 2-4 + 1 x 2-5+ 1 x 2-6 = 0.5+0+0.125+0+0.03125+0.015625 = 0.671825

Exponent (E) = 111110002 = 24810

Decimal Value = (-1)1 (1.671825) x 2248 - 127 =  

(a)

Given Floating Point number = 2

Sign bit is 1. So number is negative and hence sign of mantissa is negative.

(b)

Exponent (E) = 111110002 = 1 x 27 + 1 x 26 + 1 x 25+ 1 x 24 + 1 x 23+ 0 x 22 + 0 x 21 + 0 x 20 = 24810

(c)

Decimal Value = (-1)1 (1.671825) x 2248 - 127 =  

The sign of adjusted exponent value is 121. It is positive.